Problem3 In the short run, a firms total costs of producing 5 units of output equal $95. If it produces 6 units its total costs will increase to $117.88. See pages 491-497 a. What is the marginal cost of the 6th unit of output? b. What is the firms average total cost of producing 5 units? c. What is the firms average total cost of producing 6 units? d) How is it that the average cost is increasing? Please explain.
Answer #1

3.

The total cost of 5 unit of output=$ 95

(a) If 6th output is produced, the total cost increases to $117.88,

Marginal cost of 6th unit of output= TC of 6 unit - TC of 5 unit

   =$117.88-$95

   =$22.88

(b)

Average total cost of producing 5 unit of output= TC/Q

   =95/5

   =$19

(c)

Average total cost of producing 6 unit of output= TC/Q

   =117.88/6

   =$19.65

(d)

The average cost is increasing because the marginal cost of 6th unit of output more than the average of 5 unit of output. So when the average cost of 6 unit of output is calculated the marginal cost of 6th unit of output is also included in the total cost.